OpenStudy (anonymous):

what is the solution set of x+2/x=(x+1)/(x-3) I need help with the process and finding the answer

OpenStudy (solomonzelman):

\(\large\color{black}{ \frac{\LARGE x+2 }{\LARGE x}=\frac{\LARGE x+1 }{\LARGE x-3} }\) like this?

OpenStudy (anonymous):

yes exactly!

OpenStudy (solomonzelman):

cross multiply.

OpenStudy (anonymous):

yes but how do i get the solution set from that?

OpenStudy (solomonzelman):

you will, don't worry....

OpenStudy (solomonzelman):

The problem: \(\large\color{royalblue}{ \frac{\LARGE x+2 }{\LARGE x}=\frac{\LARGE x+1 }{\LARGE x-3} }\) The steps to take: ~ cross multiply. ~ expand on both sides to complete the cross multiplication. ~ subtract \(\large\color{black}{x^2 }\) from both sides. ~ add \(\large\color{black}{ x }\) to both sides. ~ divide both sides by \(\large\color{black}{ 2 }\).

OpenStudy (anonymous):

i got x^2+x/x^2-6x-6 from cross multiplying and im confused

OpenStudy (solomonzelman):

when you cross multiply: